Consider a political discussion group consisting of 9 Democrats, 4 Republicans, and 7 Independents. Suppose that two group members are randomly selected, in succession, to attend a political convention. Find the probability of selecting an independent and then a Democrat.
a. 63/380
b. 7/380
c. 63/400
d. 9/380

Found 2 solutions by ikleyn, math_helper:
Answer by ikleyn(52792)

.

(7/20*(9/19),

where 20 = 9 + 4 + 7 is the total population and 19 = 20-1.



Answer by math_helper(2461)


Total # of people: 9+4+7 = 20

Notice that there is a pool of 20 to choose from for the first person (independent), and once that person is picked, 19 are left to choose from for the other person (democrat):

(7/20)(9/19) = _____